# MotionCLR: Motion Generation and Training-free Editing via Understanding Attention Mechanisms

### v1

@@ 4 Versatile Applications via Attention Manipulations
Motion emphasizing and de-emphasizing. In the text-driven motion generation framework, the process is driven by the input text. As discussed in Sec. 3.3, the verb of the action will be significantly activated in the cross-attention map when the action is executed. As shown in Fig. 4(a), if we increase/decrease the attention value of a verb in the cross-attention map, the corresponding action will be emphasized/de-emphasized. Besides, this method can also be extended to the grounded motion generation, which will be introduced in Sec. 6.
Motion erasing. Motion erasing is a special case of motion de-emphasizing. We treat it as a special case of motion de-emphasizing. When the decreased (de-emphasized) cross-attention value of an action is small enough, the corresponding action will be erased.
@@
In-place motion replacement. In real scenarios, we would like to edit some local motion contents of the generated result. Assuming we generate a reference motion at first, we would like to replace one action in the reference motion with another in place. Therefore, the batch size of inference examples is two during the inference stage, where the first is the reference motion and the other is the edited motion. As discussed in Sec. 3.3, the cross-attention map determines when an action happens. Motivated by this, we replace the cross-attention map of the edited motion as the one of the reference motion. As shown in Fig. 4(b), we use the replaced attention map to multiply the value matrix (text features) to obtain the output.
@@
Motion sequence shifting. It is obvious that the generated motion is a combination of different actions along the time axis. Sometimes, users would like to shift a part of the motion along the time axis to satisfy the customized requirements. As shown in Fig. 4(c), we can shift the motion sequentiality by shifting the self-attention map. As discussed in Sec. 3.3, self-attention is only related to the motion feature without related to the semantic condition, which is our motivation on manipulating the self-attention map. Thanks to the denoising process, the final output sequence should be a natural and continuous sequence.
@@
Example-based motion generation. As defined by Li et al. (2023b), example-based motion generation aims to generate novel motions referring to an example motion. In MotionCLR system, this task is a special case of the motion sequence shifting. That is to say, we can shuffle the queries of the self-attention map to obtain the diverse motions referring to the example.
Motion style transfer. As discussed in the technical details of the self-attention mechanism, the values mainly contribute to the contents of motion and the attention map determines the selected indices of motion frames. When synthesizing two motion sequences ($\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ respectively), we only need to replace $\mathbf{Q}$s in $\mathbf{M}_{2}$ with that in $\mathbf{M}_{1}$ to achieve the style of $\mathbf{M}_{2}$ into $\mathbf{M}_{1}$’s. Specifically, queries ($\mathbf{Q}$s) in $\mathbf{M}_{2}$ determine which motion feature in $\mathbf{M}_{2}$ is the most similar to that in $\mathbf{M}_{1}$ at each timestep. Accordingly, these most similar motion features are selected to compose the edited motion. Besides, the edited motion is with the motion content of $\mathbf{M}_{2}$ while imitating the motion style of $\mathbf{M}_{1}$.
We leave more technical details and pseudo codes in Appendix F.
5 Experiments
@@
Table 1: Comparison with different methods on the HumanML3D dataset. The “” notation denotes the DDIM sampling inference design choice and the other is the DPM-solver sampling choice.
@@
(b) Visualization of the edited motions on different (de-)emphasizing weight settings.
Figure 5: Motion (de-)emphasizing. Different weights of “jump” ($\uparrow$ or $\downarrow$) in “a man jumps.”.
5.1 MotionCLR Model Evaluation
The implementations of the MotionCLR are in Sec. E.1. We first evaluate the generation performance of the MotionCLR. We extend the evaluation metrics of previous works (Guo et al., 2022a). (1) Motion quality: FID is adopted as a metric to evaluate the distributions between the generated and real motions. (2) Motion diversity: MultiModality (MModality) evaluates the diversity based on the same text and diversity calculates variance among features. (3) Text-motion matching: Following Guo et al. (2022a), we calculate the R-Precision to evaluate the text-motion matching accuracy and MM-Dist to show the distance between texts and motions. The results are shown in Tab. 1, indicating a comparable performance with the state-of-the-art method. Especially, our result has a higher text-motion alignment over baselines, owing to the explicit fine-grained cross-modality modeling. As shown in Tab. 1, both DDIM and DPM-solver sampling work consistently well compared with baselines. We leave more visualization and qualitative results in Appendix A.
5.2 Inference Only Motion Editing
weight TMR-sim. FID
- 0.60 52.059 0.776
- 0.50 52.411 0.394
- 0.40 53.294 0.235
- 0.30 53.364 0.225
baseline 53.956 0.217
+ 0.30 54.311 0.210
+ 0.40 54.496 0.208
+ 0.50 54.532 0.223
+ 0.60 54.399 0.648
Table 2: Abaltion on motion (de-)emphasizing.
[…]
@@ 5.2 Inference Only Motion Editing
(a) Root height and the motion of “a man jumps.” before editing.
@@
(c) Motion in-place replacement results of a motion including multiple actions.
Figure 7: In-place motion replacement. Replacing the “jumps” in “a man jumps.” as “walks”, the edited motion and the vanilla motion share the same temporal area of the action execution.
In-place motion replacement. Different from naïve replacing prompts for motion replacement, in-place motion replacement not only replaces the original motion at the semantic level, but also needs to replace motions at the exact temporal place. Fig. 7(a) and Fig. 7(b) show the root height trajectory and the root horizontal velocity, respectively. In this case, the edited and original motion share the same time zone to execute the action. Besides, the edited motion is semantically aligned with the “walk”. Fig. 7(c) also shows results of replacing “runs” as“jumps” without changing the sitting action.
Motion sequence shifting. Here, we provide some comparisons between the original motion and the edited one. In Fig. 9, we take “ ” and “ ” to represent different time bars, whose orders represent the sequentially. As can be seen in Fig. 9(a), the execution of waving hands is shifted to the beginning of the motion. Besides, as shown in Fig. 9(b), the squatting action has been moved to the end of the motion. These results show that the editing of the self-attention map sequentiality has an explicit correspondence with the editing motion sequentially. More results are in Sec. A.6.
Example-based motion generation. The example-based motion generation (Li et al., 2023b) task has two basic requirements. (1) The first one is the generated motions should share similar motion textures (Li et al., 2002) with the example motion. We observe the high-dimension structure of motions via dimensionality reduction. As the t-SNE visualization results shown in Fig. 8, generated motions driven by the same example are similar to the given example (in the same color). (2) Besides, different generated motions driven by the same example should be diverse. As shown in Fig. 10, these generated results are diverse not only in local motions (Fig. 10(a)) but also in the global trajectory (Fig. 10(b)). Furthermore, results in Fig. 10 also share the similar motion textures. We leave more visualization results in Sec. A.7.
@@
Figure 9: Comparison between original and shifted motions. Time bars are shown in different colors. (a) The original figure raises hands after the walking action. The shifted one has the opposite sequentiality. (b) The key squatting action is shifted to the end of the sequence, and the standing-by action is shifted to the beginning.
@@
Figure 10: Diverse generated motions driven by the same example. Prompt: “a person steps sideways to the left and then sideways to the right.”. (a) The diverse generated motions driven by the same example motion share similar movement content. (b) The root trajectories of diverse motions are with similar global trajectories, but not the same.
@@
(b) Style reference: a man is doing hip-hop dance”, Content reference: a person runs around a circle”. The stylized result shows a running motion with bent hands, shaking left and right.
@@
Motion style transfer. As shown in Fig. 11, in the MotionCLR framework, the style reference motion provides style and the content reference motion provides keys and values. As can be seen in Fig. 11, all edited results are well-stylized with style motions and keep the main movement content with the content reference.

### v2

@@ 4 Motion Editing Applications via Attention Manipulations
4.1 In-place Motion Replacement
@@
In real scenarios, we would like to edit some local motion contents of the generated result. The key challenge is to keep the original composition and replace it with the new motion contents. Assuming we generate a reference motion at first, we would like to replace one action in the reference motion with another in place. Therefore, the batch size of inference examples is two during the inference stage, where the first is the reference motion and the other is the edited motion. As discussed in Sec. 3.3, the cross-attention map determines when an action happens. Motivated by this, we replace the cross-attention map of the edited motion as the one of the reference motion. As shown in Fig. 4(a), we use the replaced attention map to multiply the value matrix (text features) to obtain the output. As a result, we will obtain the motion of edited semantics with referenced temporal composition.
4.2 Motion Emphasis and De-emphasis
In the text-driven motion generation framework, the process is driven by the input text. As discussed in Sec. 3.3, the verb of the action will be significantly activated in the cross-attention map when the action is executed. As shown in Fig. 4(b), if we increase/decrease the attention value of a verb in the cross-attention map, the corresponding action will be emphasized/de-emphasized, which can be implemented by multiplication ($\mathbf{A}_{:,i}\leftarrow\mathbf{A}_{:,i}\times(1+\alpha)$). Here, positive and negative values of $\alpha$ represent the motion emphasis and de-emphasis, respectively. Besides, this method can also be extended to the grounded motion generation application, which will be introduced in Sec. 6.
4.3 Motion Erasing
Motion erasing is a special case of motion de-emphasis. We treat it as a special case of motion de-emphasis. When the decreased (de-emphasized) cross-attention value of an action is small enough, the corresponding action will be erased. The difference with motion de-emphasis is that the motion erasing is applied in a sub-temporal area specified by users in the whole sequence, and the de-emphasis is applied in the whole sequence.
Despite these semantic editing applications, MotionCLR also supports other editing tasks by manipulating self-attention.
4.4 Motion Sequence Shifting
@@
It is obvious that the generated motion is a combination of different actions along the time axis. Sometimes, users would like to shift a part of the motion along the time axis to satisfy the customized requirements. As shown in Fig. 4(c), we can shift the motion sequentiality by shifting the self-attention map. As discussed in Sec. 3.3, self-attention is only related to the motion feature without related to the semantic condition, which is our motivation on manipulating the self-attention map. Thanks to the denoising process, the final output sequence should be a natural and continuous sequence.
4.5 Example-based Motion Generation
@@
As defined by Li et al. (2023b, 2002), example-based motion generation aims to generate novel motions referring to an example motion. In MotionCLR system, this task is a special case of the motion sequence shifting. That is to say, we can shuffle the self-attention map along the temporal axis to obtain the diverse motions referring to the example. As analyzed in Sec. 3.3, the “value”s ($\mathbf{V}$) in self-attention means the texture of the motion. Therefore, shuffling the self-attention map along the temporal axis without manipulating “value”s compromises the similar motion texture to the original one.
Table 1: Comparison with different methods on the HumanML3D dataset. The baselines include diffusion-based methods and state-of-the-art methods. The “” notation denotes the DPM-solver sampling inference design choice and “” is the DDIM sampling choice. As DPM-solver and DDIM present comparable performance, without specification, we set the DDIM sampling as our default choice. The comparison shows that MotionCLR is with comparable performance with state-of-the-art methods.
@@
4.6 Motion Style Transfer
As discussed in the technical details of the self-attention mechanism, the values mainly contribute to the contents of motion and the attention map determines the selected indices of motion frames. Following Raab et al. (2024a), when synthesizing two motion sequences ($\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ respectively), we only need to replace $\mathbf{Q}$s in $\mathbf{M}_{2}$ with that in $\mathbf{M}_{1}$ to achieve the style of $\mathbf{M}_{2}$ into $\mathbf{M}_{1}$’s. Specifically, queries ($\mathbf{Q}$s) in $\mathbf{M}_{2}$ determine which motion feature in $\mathbf{M}_{2}$ is the most similar to that in $\mathbf{M}_{1}$ at each timestep. Accordingly, these most similar motion features are selected to compose the edited motion. Besides, the edited motion is with the motion content of $\mathbf{M}_{2}$ while imitating the motion style of $\mathbf{M}_{1}$.
5 Experiments
5.1 Implementation Details
The MotionCLR model is trained on the HumanML3D dataset with one NVIDIA A-100 GPU based on PyTorch (Paszke et al., 2019). The latent dimension of the motion embedding is $512$. We take the CLIP-ViT-B model to encoder text as word-level embeddings. The training process utilizes a batch size of 64, with a learning rate initialized at $42e-$ 4 and decaying at a rate of $0.9$ every $5,000$ steps. Additionally, a weight decay of $21e-$ 2 is employed to regularize the model parameters. For the diffusion process, the model is trained over $1,000$ diffusion steps. We incorporate a probability of $0.1$ for condition masking to facilitate classifier-free guidance learning. During training, dropout is set at $0.1$ to prevent overfitting, and all networks in the architecture follow an $8$-layer Transformer design. For motion representation, we follow the setting in Guo et al. (2022a).
[…]
5.2 Motion Generation Evaluation for MotionCLR
The implementation details of the MotionCLR are in Sec. 5.1. We first evaluate the generation performance of the MotionCLR on HumanML3D Guo et al. (2022a). We extend the evaluation metrics of previous works (Guo et al., 2022a), including FID, R-Precision, MM-Dist, and, Muiti-Modality. The results are shown in Tab. 1, indicating a comparable performance with the state-of-the-art method. Especially, our result has a higher text-motion alignment over baselines, owing to the explicit fine-grained cross-modality modeling. As shown in Tab. 1, both DDIM and DPM-solver sampling work consistently well compared with baselines. We leave more visualization and qualitative results in Appendix A.
5.3 Constructing Evaluation Set for Motion Editing
To evaluate the semantic correspondence between cross-attention and the generated motion, we construct an evaluation set to verify the observation qualitatively. We label verbs in HumanML3D texts in each sentence. Here, only a single verb in the sentence will be labeled and the sentences without any verbs will be filtered out. As a result, we construct 19,492 texts for evaluation, namely the HVerb test set. For the convenience of validating the in-placement replacement task, we also assign a new verb as the replacing verb of the original verb in the HVerb test set. To additionally test the explainability of the model in-the-wild, we construct 200 text prompts with verb labels via GPT-4o Achiam et al. (2023), namely HVerb-wild, whose results are checked by researchers.
Table 2: IoU (%) metrics on different settings. High coherence among E1, E2, and E3 shows the fine-grained text-motion modeling in cross attention.
Experiment ID E1 E2 E3 E4 E5
HVerb 74.3% 72.9% 77.8% 17.5% 17.0%
Experiment ID E1 E2 E3 E4 E5
HVerb-wild 73.5% 71.4% 74.5% 18.5% 19.4%
5.4 Quantitative Results for Correspondence between Cross Attention and Generated Motion
Before evaluating the motion editing result, we initially validate the word-motion correspondence in the cross-attention map of MotionCLR. We set up 5 groups of experiments for comparison. (1) E1: IoU between cross attention & root velocity. We treat the value in the cross-attention map larger than 80% of the maximum value as an activated action. Accordingly, we treat the activated and unactivated parts as 1 and 0 respectively. Similarly, we also apply this to the root velocity to get another activation map. The IoU metric between these two activation maps means the coherence between the attention activation and the kinematic features. (2) E2: IoU between the cross attention & moment retrieval value. We use the moment retrieval function of TMR to calculate the verb-motion similarity in 20-frame windows, obtaining the similarity between sub-motion and the verb along the whole motion sequence. Accordingly, we can also calculate a metric between attention and the moment retrieval value (also set 60% as the threshold). (3) E3: IoU between the root velocity & moment retrieval value. This setting is similar to that in E2 and E3. Both root velocity and moment retrieval value are explicit kinematic metrics used to describe action execution. The motivation to set E3 is for setting a comparison group to verify that the cross-attention correspondence is coherent with these two metrics. (4): E4: neg. cross-attention & root velocity. We modify the setting in E1 and replace the cross-attention map with a randomly sampled cross-attention map in the test in calculating IoU with the root velocity. (5): E5: neg. cross-attention & moment-retrieval-value. Similarly, we modify the setting in E2 and replace the cross-attention map with a randomly sampled cross-attention map in the test in calculating IoU with moment-retrieval-value.
As can be seen in Tab. 2, compared with E3 (both explicit action indicators: speed, moment retrieval of TMR), the values in E1/E2 are similar to those in E3. Therefore, the cross-attention activation is well aligned with motion execution. Besides, as shown in random sampling comparison groups (E1 v.s. E4/E5), the cross attention is aligned with the action execution in the motion in the generation process of each motion. The result proves the word-motion correspondence in MotionCLR.
5.5 Evaluation on Inference-only Motion Editing
Table 3: In-place motion replacement. The “R” and “C” settings represent “Replace A as B” and “Change A with B” prompts of MotionFix. The light gray text is the comparison group, denoting the upper bound of the performance. The bold numbers are the best results excepting the comparison group. The significant metric margin over baselines shows the good performance of the method, even some methods requiring specific training.
HVerb
training-
free
align with
original text
(%) $\downarrow$
align with
edited text
(%) $\uparrow$
unedited part
preserving
(mm) $\downarrow$
w/o editing - 76.205 64.687 0.0
editing text only 62.324 68.368 201.5
MotionFix (“R”) 74.526 63.542 130.2
MotionFix (“C”) 75.112 63.780 142.1
Ours 63.324 68.125 57.9
HVerb-wild
training-
free
align with
original text
(%) $\downarrow$
align with
edited text
(%) $\uparrow$
unedited part
preserving
(mm) $\downarrow$
w/o editing - 73.678 59.229 0.0
editing text only 56.235 66.701 235.0
MotionFix (“R”) 70.125 58.560 151.1
MotionFix (“C”) 71.588 60.009 149.0
Ours 58.124 65.976 59.8
(a) Root height visualization and the motion of “a man jumps.” before applying editing.
[…]
(c) Motion in-place replacement results of a motion including multiple actions. In this case, we replace the “jumps” in “a man jumps.” as “walks” without editing the sitting action. The edited motion and the vanilla motion share the same temporal area of the action execution.
Figure 5: In-place motion replacement. (a) and (b) are a pair of motions before and after editing. (c) is a comparison of original and edited motions.
5.5.1 In-place motion replacement.
Different from naïve replacing prompts for motion replacement, in-place motion replacement not only needs to replace the original motion at the semantic level, but also needs to replace motions at the exact temporal place. Fig. 5(a) and Fig. 5(b) show the root height trajectory and the root horizontal velocity, respectively. In this case, the edited and original motion share the same time zone to execute the action. Besides, the edited motion is semantically aligned with the word “walk”. Fig. 5(c) also shows results of replacing “runs” as“jumps” without changing the sitting action. These quantitative results show the effectiveness of the in-place motion replacement application.
Figure 6: Results of editing multiple semantics from the same motion. All edited motions with different semantics share the same temporal location as the original motion.
For the in-place motion replacement application, we additionally provide replacement results in Fig. 6. As shown in Fig. 6, MotionCLR changes the original “walk” motion into “jumps” and “dances” at the corresponding place, serving as a diverse semantic editing function. All edited motion with different semantics share the same action temporal location with the original motion. This function shows the robustness of the method for diverse editing requirements.
We also test our method on the HVerb and the HVerb-wild test sets quantitatively, which are shown in Tab. 3. We take the latest motion editing method, MotionFix, as a baseline, whose prompts are set as ‘Replace A as B” or “Change A with B”. For example, the prompt can be “Replace jump as walk”. Two examples are set to enhance prompt diversity of the baseline. As the editing process of COMO Huang et al. (2025) and Goel et al. (2024) are not fully open-sourced before writing this paper, we do not take these as our baselines. For the application of in-place motion replacement, the editing goal comes up with two aspects. (1) The semantics of the edited motion should be aligned with the replaced text. Here, we take the TMR Petrovich et al. (2023) similarity (0$\sim$100%) to evaluate the text-motion similarity. As shown in Tab. 3, the former two columns indicate that the edited motion is more semantically aligned with the edited text, and less aligned with the original text. (2) The unedited part of the motion should be reserved. We use the TMR moment retrieval function and the annotated verb to filter out the editing area of the motion, similar to Sec. 5.4. For the filtered motion part, we take the MPJPE (mm) metric to evaluate the motion-preserving ability of the unedited motion. As can be seen in the last column of Tab. 3, MotionCLR shows a stable motion-preserving ability on unedited areas. As the dataset construction process of MotionFix is based on a similar motion retrieval process, the semantic changes in the editing process are unsatisfactory.
@@ 5.5 Evaluation on Inference-only Motion Editing
(b) Motion visualization of the edited motions on different (de-)emphasis weight settings.
Figure 7: Motion (de-)emphasis. Different weights of “jump” ($\uparrow$ or $\downarrow$) in “a man jumps.”.
@@
5.5.2 Motion (de-)emphasis and motion erasing.
@@
Additionally, we would like to discuss the difference between reweighting the cross-attention map and adjusting classifier-free guidance weights. For intuitive understanding, the classifier-free guidance weight also controls the semantics of the motion with the input text. However, as the classifier-free guidance mainly works for the semantic alignment between text and motion, it cannot control the weight of each word. We take the sentence “a man jumps.” as an example for a fair comparison, which is the case used in the main text (suggested to refer to Fig. 7(a) for comparison). As shown in Fig. 9, the generated motions with different $w$ values illustrate that $w$ cannot influences both the height and frequency of the jump. Nevertheless, the classifier-free guidance is limited in its ability to control more detailed aspects, such as the exact height and number of actions. Therefore, while $w$ improves text-motion alignment, it cannot achieve fine-grained adjustments. As there is no benchmark for such an application, we quantitatively evaluate this in the user study part (Sec. 5.7).
As motion erasing is a special case of motion de-emphasis, we do not provide more quantitative on this application. We provide some visualization results in Fig. 8. As can be seen in Fig. 8(a), the second jumping action is erased. Besides, the “waving hand” case shown in Fig. 8(b) shows that the final 1/3 waving action is also removed.
@@
Figure 10: Comparison between original and shifted motions. Time bars are shown in different colors. (a) The original figure raises hands after the walking action. The shifted one has the opposite sequentiality. (b) The squatting action is shifted to the end of the sequence, and the standing-by action is shifted to the beginning.
5.5.3 Motion sequence shifting.
Here, we provide some comparisons between the original motion and the edited one. In Fig. 10, we take “ ” and “ ” to represent different time bars, whose orders represent the sequentially. As can be seen in Fig. 10(a), the execution of waving hands is shifted to the beginning of the motion. Besides, as shown in Fig. 10(b), the squatting action has been moved to the end of the motion. These results show that the editing of the self-attention map sequentiality has an explicit correspondence with the editing motion sequentially. More results are in Sec. A.3.
@@
Figure 11: Diverse generated motions driven by the same example. Prompt: “a person steps sideways to the left and then sideways to the right.”. (a) The diverse generated motions driven by the same example motion share similar movement content. (b) The root trajectories of diverse motions are with similar global trajectories but not the same.
5.5.4 Example-based motion generation.
The example-based motion generation (Li et al., 2023b) task has two basic requirements. (1) The first one is the generated motions should share similar motion textures (Li et al., 2002) with the example motion. We observe the high-dimensional structure of motions via dimensionality reduction. As the t-SNE visualization results shown in Fig. 12, generated motions driven by the same example are similar to the given example (in the same color). (2) Besides, different generated motions driven by the same example should be diverse. As shown in Fig. 11, these generated results are diverse not only in local motions (Fig. 11(a)) but also in the global trajectory (Fig. 11(b)). Furthermore, results in Fig. 11 also share the similar motion textures. We leave more visualization results in Sec. A.4.
[…]
Table 4: Comparison with baselines. Our method enjoys better editing quality and higher diversity than previous methods. The motion coverage is comparable with the state-of-the-art methods, although our method is not designed for the specific task.
specific FID $\downarrow$ Diversity $\uparrow$ Coverage (%) $\uparrow$
GANimator (2022) 0.505 2.012 84.1
GenMM (2023b) 0.514 2.478 97.5
MotionCLR 0.427 2.567 97.0
Table 5: Comparison on FID and diversity values with manipulating self-attention in the motion space of the denoising process. As can be seen in the table, our method enjoys better editing quality and higher diversity than editing at each diffusion step.
FID $\downarrow$ Diversity $\uparrow$
Diff. manipulation 0.718 1.502
MotionCLR manipulation 0.427 2.567
As the diffusion denoising process can manipulate the motion directly in the denoising process, we provide a baseline for comparison with our motion shifting and example-based motion generation applications. Here, for convenience, we only take the example-based motion generation application as an example for discussion. In this section, we conduct a comparison between our proposed editing method and diffusion manipulation in the motion space, focusing on the FID and diversity metrics. The 200 samples used in this experiment were constructed by researchers. As depicted in Tab. 5, the “Diff. manipulation” serves for our comparison. Our method achieves an FID value of 0.427, indicating a relatively high generation quality, while the “Diff. manipulation” achieves a higher FID of 0.718, demonstrating worse fidelity. Conversely, in terms of diversity, the “MotionCLR manipulation” exhibits a higher diversity score of 2.567 compared to the 1.502 of the “Diff. manipulation.” These results verify our method is better than manipulating noisy motions in the denoising process. The main reason for the better quality and diversity mainly relies on the many times of manipulation of self-attention, but not the motion. Directly manipulating the motion results in some jitters, making more effort for models to smooth. Besides, the shuffling times of manipulating the self-attention maps are higher than the baseline, contributing to the better diversity.
@@
(b) Style reference: “a man is doing hip-hop dance”, Content reference: “a person runs around a circle”. The stylized result shows a running motion with bent hands, shaking left and right.
@@
Table 6: Comparison of motion style transfer across baselines. The “Gen.” and “Inv.” settings represent the editing during generation and DDIM inversion Song et al. (2021) (following Raab et al. (2024a)) settings.
FID $\downarrow$
R Precision
(Top 3)
$\uparrow$
Texture Ref. Foot
Contact Sim.
$\uparrow$
Style Ref.
Loc.Sim.
$\uparrow$
AITS $\downarrow$
MoMo (Gen.) (2024a) 2.33 0.439 0.816 0.972 0.544
MoMo (Inv.) (2024a) 2.50 0.490 0.793 0.856 0.691
Ours (Gen.) 0.65 0.749 0.877 0.989 0.345
Ours (Inv.) 0.69 0.784 0.851 0.914 0.432
5.5.5 Motion style transfer.
As shown in Fig. 13, in the MotionCLR framework, the style reference motion provides style and the content reference motion provides keys and values. As can be seen in Fig. 13, all edited results are well-stylized with style motions and keep the main movement content with the content reference. To qualitatively evaluate the transferred result, we compare our method with the latest method, MoMo Raab et al. (2024a). We follow the evaluation protocol in Raab et al. (2024a), where the comparison results in Tab. 6 indicate the better transfer result than MoMo. The quality gain mainly comes from the better base model of MotionCLR and the disentangled modeling of self-attention and cross-attention. Besides, we also compare the speed (AIST metric following Chen et al. (2023)) with baselines, indicating better efficiency of MotionCLR.